OpenStudy (anonymous):

how many 3 digit natural numbers are divisible by 4 ( medal to the correct answer)

OpenStudy (anonymous):

Three digit natural numbers are from 100 to 999.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got the answer as 223 but my textbook says it is 225

OpenStudy (anonymous):

996=249*4 100=25*4 249-25+1=225

OpenStudy (anonymous):

Wait i will tell u all how i did it

ganeshie8 (ganeshie8):

oh you wanto divide by 4

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

100 -> 999 100 + 4k k = 0, 1, ....

ganeshie8 (ganeshie8):

100+4k < 1000 solve k

OpenStudy (anonymous):

1st term ---- 100 2nd term ----104 3rd term ----108 4th term ----112 We know tn= a + (n-1)d nth term in this arithmetic progression would be 996 i suppose So, 996 =100+(n-1)4 It comes to 223

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

looks good, except for the arithmetic error in ur last line

ganeshie8 (ganeshie8):

996 =100+(n-1)4 subtract 100 both sides 896 = (n-1)4 divide 4 both sides 224 = n-1 add 1 both sides 225 = n